# Think Before You Talk: Enhancing Meaningful Dialogue Generation in Full-Duplex Speech Language Models with Planning-Inspired Text Guidance

### v1

@@ Abstract
Full-Duplex Speech Language Models (FD-SLMs) are specialized foundation models designed to enable natural, real-time spoken interactions by modeling complex conversational dynamics such as interruptions, backchannels, and overlapping speech, and End-to-end (e2e) FD-SLMs leverage real-world double-channel conversational data to capture nuanced two-speaker dialogue patterns for human-like interactions. However, they face a critical challenge—their conversational abilities often degrade compared to pure-text conversation due to prolonged speech sequences and limited high-quality spoken dialogue data. While text-guided speech generation could mitigate these issues, it suffers from timing and length issues when integrating textual guidance into double-channel audio streams, disrupting the precise time alignment essential for natural interactions. To address these challenges, we propose TurnGuide, a novel planning-inspired approach that mimics human conversational planning by dynamically segmenting assistant speech into dialogue turns and generating turn-level text guidance before speech output, which effectively resolves both insertion timing and length challenges. Extensive experiments demonstrate our approach significantly improves e2e FD-SLMs’ conversational abilities, enabling them to generate semantically meaningful and coherent speech while maintaining natural conversational flow. [^1] Demos and code will be available at https://dreamtheater123.github.io/TurnGuide-Demo/ and https://github.com/dreamtheater123/TurnGuide.
@@ 1 Introduction
Speech Language Models (SLMs) have recently attracted significant attention for their role as foundation models in facilitating speech-based interactions with human users (Cui et al. 2024; Hurst et al. 2024). These models leverage advanced autoregressive modeling techniques to produce coherent speech aligned with contextual cues. To facilitate seamless and natural communication experiences, Full-Duplex Speech Language Models (FD-SLMs), a specialized type of SLM, aspire to emulate human-like conversations (Nguyen et al. 2023; Wang et al. 2025). FD-SLMs model user channel and assistant channel audio data simultaneously to support real-time communication characterized by intricate spoken conversational dynamics, including interruptions (Chen et al. 2025), backchannels (Défossez et al. 2024), overlapping speech (Nguyen et al. 2023), and other features that closely mimic natural human interactions. Existing FD-SLMs can be categorized as cascaded and end-to-end (e2e) systems. Cascaded FD-SLMs utilize a separate prediction module to classify spoken dialogues into predefined states, guiding SLM generation (Chen et al. 2025; Wang et al. 2024). While demonstrating efficacy in handling elementary duplex features like interruptions, they are inherently bounded by their dependence on explicit rule-based frameworks. In contrast, e2e FD-SLMs, which we adopt, leverage real-world conversational data to train models end-to-end, capturing nuanced dialogue patterns for more natural, human-like interactions (Wang et al. 2025; Veluri et al. 2024). Figure 1 illustrates these two FD-SLM types. The right side depicts the complex conversational patterns in natural human dialogues, which are modeled by e2e FD-SLMs.
@@
While e2e FD-SLMs facilitate more natural conversational experiences, a significant challenge arises when training these models on spoken dialogue data. Their conversational abilities tend to degrade in comparison to text-based Large Language Models (TLMs) due to the nature of prolonged speech sequences and the scarcity of high-quality speech training data (Zeng et al. 2024; Xie and Wu 2024). One promising solution to this challenge is to transfer the text abilities to speech abilities by leveraging text-guided speech generation, where the model first generates the assistant’s text response based on the input context, followed by generating the corresponding speech response. However, applying text guidance in e2e FD-SLMs poses significant challenges, as these models require precise time alignment between the user and assistant audio streams to effectively capture complex conversational patterns and provide timely, meaningful responses. Integrating text elements into the audio streams can easily disrupt this alignment, ultimately compromising the responsiveness and coherence of the generated speech.
@@
The challenges of integrating texts into assistant-channel audio data can be summarized into two main aspects. 1) Insertion Timing. The text must be inserted precisely at the moment the assistant begins speaking. If the insertion is too early, the assistant may respond without accessing the full context of the user’s speech, leading to misinterpretations or hallucinations. Conversely, if the insertion is too late, the assistant’s speech may precede the inserted text, rendering the text guidance approach ineffective. 2) Insertion Length. The length of the inserted text is critical. If the text is too lengthy, the assistant may continue speaking excessively and struggle to adapt its responses to subsequent user inputs. On the other hand, if the inserted text is too brief, the semantic content may become overly fragmented, diminishing the effectiveness of the text guidance approach (Yao et al. 2025).
@@
To address these challenges, we propose a novel approach, TurnGuide, to seamlessly integrate text data into real-world double-channel spoken conversational data for the training of e2e FD-SLMs. Our approach draws inspiration from the human planning process during conversations, where individuals think ahead, assess real-time speech by others, and plan what to say, when to say it, and for how long. Specifically, we dynamically segment the assistant’s speech into turns, which represent the reasonable lengths of spoken content that convey the assistant’s intended expression, reflecting the structured planning process behind conversational flow. The assistant is trained to plan and output text guidance for each turn before outputting the corresponding speech. By doing so, our method effectively addresses both insertion timing and insertion length challenges. Extensive experiments demonstrate that our approach significantly improves e2e FD-SLMs’ conversational ability, enabling them to generate semantically meaningful and coherent speech while maintaining the natural conversational flow of the generated full duplex dialogue. To summarize our contributions:
  1. 1.
    We propose a novel planning-inspired approach, TurnGuide, which plans both the textual content and the turn-taking timing before speaking, enabling the effective text-guided speech generation for e2e FD-SLMs.
  2. 2.
@@
    Our approach includes a dynamic segmentation and text-guided dialogue modeling framework, addressing the core challenges of insertion timing and length in double-channel dialogues.
  3. 3.
    We demonstrate through extensive experiments that our proposed approach greatly enhances the capability of FD-SLMs to produce semantically meaningful and coherent speech, while preserving the natural conversational flow in full-duplex spoken dialogues.
@@ 2 Related Works
Figure 2: Illustration of the multi-modal turn segmentation and alignment framework, which dynamically segments the dialogue speech into turns and aligns the text with speech turns. The text turns serve as the contents derived by the planning process.
@@ 2.1 Full Duplex Speech Language Models
Full Duplex Speech Language Models (FD-SLMs) process double-channel conversational data to interact with users seamlessly. FD-SLMs can be categorized into cascaded and end-to-end (e2e) approaches.
Cascaded FD-SLMs
@@
The main focus of cascaded FD-SLMs is on defining states. For instance, LSLM (Ma et al. 2025) uses states like ”interrupted” and ”keep speaking”, where interruptions are triggered by specific keywords. Mini-Omni 2 (Xie and Wu 2024) specifies ”start speaking” and ”stop speaking,” halting generation upon hearing ”Stop Omni!”. Freeze-Omni (Wang et al. 2024) bases states on whether the user has finished speaking, prompting the model to wait or interrupt. Minmo (Chen et al. 2025) simplifies its states to whether the model should speak up or keep listening. All these methods rely on synthetic conversational data to manage state transitions, limiting their ability to foster natural, human-like spoken interactions.
End-to-end FD-SLMs
E2e FD-SLMs, on the other hand, employ a transformer-based foundation model architecture (Vaswani et al. 2017; Achiam et al. 2023) to learn from real-world double-channel conversational audio data. This e2e approach enables SLMs to autonomously capture intricate conversational dynamics from authentic human interactions, enabling more natural conversations.
The main focus of e2e FD-SLMs is on effectively modeling double-channel audio data and capturing the interactions between these channels. dGSLM (Nguyen et al. 2023) introduces a Dialogue Language Model that employs distinct transformers to process each audio channel separately, using cross-attention mechanisms to model their interactions. SyncLLM (Veluri et al. 2024) takes a different approach by jointly modeling the double-channel audio within a single LLM, interleaving the two audio sequences into a unified sequence. Moshi (Défossez et al. 2024) adopts an RQ-Transformer architecture (Lee et al. 2022), designed to predict the assistant channel’s output based on the user channel’s input at each time step. NTPP (Wang et al. 2025) proposes a Next-Token-Pair Prediction training objective, enabling simultaneous prediction of audio tokens for both channels in each step.
@@ 2.2 Text-Guided Speech Generation in SLMs
SLMs enable end-to-end spoken interactions by generating output speech based on user speech input. However, studies have shown that SLMs exhibit significantly lower performance compared to TLMs counterparts (Cui et al. 2025; Chen et al. 2025; Long et al. 2025). To mitigate this, researchers focus on integrating text to guide the speech generation process, aiming to enhance SLMs by transferring capabilities from TLMs.
Transferring abilities from TLMs to SLMs involves two key processes: text-speech alignment and end-to-end speech interaction. Text-speech alignment ensures the connection between text and speech representations within the SLM backbone, enabling knowledge transfer from text modality to speech modality. This alignment is commonly achieved through Automatic Speech Recognition (ASR) and Text-To-Speech (TTS) tasks during SLM pre-training (Zeng et al. 2024; Wang et al. 2024; Chen et al. 2025). Some approaches also focus on partially replacing speech data with its corresponding text, creating interleaved text-speech sequences for pre-training purposes (Nguyen et al. 2025; Zeng et al. 2024). Following this, SLMs are trained using speech interaction datasets to achieve seamless end-to-end capabilities for speech-based interactions, effectively enhancing their performance and functionality.
The exploration of the text-guided speech generation for e2e FD-SLMs remains in its infancy. Currently, the only notable study in this area is Moshi (Défossez et al. 2024), which proposes adding text tokens corresponding to their respective speech sounds as they occur. However, this method fragments text semantics excessively and does not significantly improve SLM performance. In contrast, our approach adopts a more holistic strategy, planning the overall context before the text insertion process. By systematically inserting text by turns, TurnGuide achieves substantial enhancements in SLM performance.
@@ 3 Methodology
Figure 3: Illustration of our text-guided full duplex dialogue modeling framework with two interleaving strategies.
In this section, we present the comprehensive methodology of TurnGuide, designed to enhance the e2e FD-SLM by utilizing text guidance. This approach is divided into two key frameworks: one for planning and segmenting spoken dialogue into turns and performing speech-text alignment, and the other for modeling text-guided full-duplex dialogue.
3.1 Multi-Modal Turn Segmentation and Alignment Framework
@@
We leverage both speech and text as multi-modal cues to perform turn segmentation, as shown in Figure 2. We first apply Voice Activity Detection (VAD) to identify speech segments within the audio stream. Let $A$ represent the input audio sequence, VAD produces a list of speech segments:
@@
where each tuple $(\tau_{i}^{start},\tau_{i}^{end})$ represents the start and end timestamps of the $i$-th detected speech segment. We use the pyannote library (Bredin et al. 2020) as the VAD module. Considering the VAD result is too fragmented, we merge nearby speech segments into Inter-Pausal Units (IPUs) (Nguyen et al. 2023), which are defined by consecutive segments having pauses within a threshold $\tau_{IPU}=0.5$ seconds, i.e.,
@@
To further refine the turn segmentation, we utilize text transcripts with word-level timestamps aligned to the speech audio. While some speech dialogue datasets include transcriptions, we use an Automatic Speech Recognition (ASR) module with timestamp functionality to perform the task. This approach ensures adaptability to unlabeled dialogue datasets. The ASR process is formulated as
@@
where $(\delta_{i}^{start},\delta_{i}^{end})$ represents the start and end timestamps of word $w_{i}$. In practice, we employ the Whisper medium model (Radford et al. 2023) for both ASR and timestamp extraction tasks. To enhance the stability of the timestamp extraction process, we utilize the whisper-timestamped package (Louradour 2023).
@@
Finally, we utilize timestamp information to enhance turn segmentation further. First, we group each word into its associated IPU using $\tau^{start}$. We incorporate a tolerance parameter $\tau_{tol}=0.6$ seconds to adjust the word segmentation boundary slightly. This adjustment accounts for instances where, in practice, the first word of an IPU might incorrectly align with the preceding IPU due to inconsistencies in timestamping. The process is defined as
$W_{j}=\{w_{i}:\tau_{IPU_{j}}^{start}-\tau_{tol}\leq\delta_{i}^{start}\leq\tau_{IPU_{j}}^{end}-\tau_{tol}\},$ (5)
@@
where $W_{j}$ represents words grouped into turn $j$, and $\tau_{IPU_{j}}^{start}$ and $\tau_{IPU_{j}}^{end}$ represent the start and end times of IPU $j$. We denote the speech segment for IPU $j$ as $S_{j}$. Then, we split both the grouped text segments and their corresponding speech segments at natural punctuation boundaries to create coherent dialogue turns. Each word group $W_{j}$ is segmented at punctuation marks $P=\{$,”, “.”, “??”, “!!}\}, i.e., $W_{j}^{final}=\text{split}(W_{j},P)$, and we extract the corresponding speech segment $S_{j}^{final}$ according to the word-level timestamps. The final output consists of aligned text-speech pairs $\{(W_{j}^{final},S_{j}^{final})\}$ where each text segment is temporally aligned with its corresponding speech segment, serving as training data for text-guided full-duplex dialogue modeling.
3.2 Text-Guided Full Duplex Dialogue Modeling

### v2

# TurnGuide: Enhancing Meaningful Full Duplex Spoken Interactions via Dynamic Turn-Level Text-Speech Interleaving

@@ Abstract
Full-Duplex Speech Language Models (FD-SLMs) are specialized foundation models designed to enable natural, real-time spoken interactions by modeling complex conversational turn-taking such as interruptions, backchannels, and overlapping speech. End-to-end (e2e) FD-SLMs leverage real-world double-channel conversational data to capture nuanced two-speaker dialogue patterns for human-like interactions, but their conversational abilities often degrade compared to pure-text conversation due to prolonged speech sequences and limited high-quality spoken dialogue data. Although interleaved text-speech generation could mitigate this degradation, integrating discrete text tokens into continuous double-channel audio streams could disrupt the precise time alignment required for fluid interaction. To address this, we propose TurnGuide, a novel text-speech interleaved generation approach for e2e FD-SLMs that dynamically segments assistant speech into dialogue turns and interleaves turn-level text and speech generation. This approach allows FD-SLMs to integrate the semantic intelligence of LLMs without compromising the natural acoustic flow. Extensive experiments show that TurnGuide not only significantly improves e2e FD-SLMs to produce semantically meaningful, coherent speech but also achieves state-of-the-art performance on various turn-taking events. [^1] Demos and code will be available at https://dreamtheater123.github.io/TurnGuide-Demo/ and https://github.com/dreamtheater123/TurnGuide.
TurnGuide: Enhancing Meaningful Full Duplex Spoken Interactions via Dynamic Turn-Level Text-Speech Interleaving
Wenqian Cui1thanks: Project lead. Email: wenqian.cui@link.cuhk.edu.hk, Lei Zhu2, Xiaohui Li2thanks: Corresponding author., Zhihan Guo1, Haoli Bai2, Lu Hou2, Irwin King122footnotemark: 2 1The Chinese University of Hong Kong, 2Huawei Technologies
@@ 1 Introduction
Speech Language Models (SLMs) have attracted attention as foundation models for speech-based interactions with human users Cui et al. (2024); Hurst et al. (2024). Full-Duplex Speech Language Models (FD-SLMs), a specialized type of SLM, model user- and assistant-channel audio data simultaneously to support real-time communication with various natural turn-taking events such as interruptions Chen et al. (2025), backchannels Défossez et al. (2024), and overlapping speech Nguyen et al. (2023), aiming to emulate human-like conversations Nguyen et al. (2023); Wang et al. (2025). Existing FD-SLMs can be categorized as cascaded and end-to-end (e2e) systems. Cascaded FD-SLMs utilize a separate prediction module to classify spoken dialogues into predefined states, guiding SLM generation Chen et al. (2025); Wang et al. (2024). While demonstrating efficacy in handling elementary turn-taking like interruptions, they are inherently bound by their dependence on explicit rule-based frameworks. In contrast, e2e FD-SLMs, which we adopt, leverage real-world conversational data to train models e2e, capturing nuanced dialogue patterns for more natural, human-like interactions Wang et al. (2025); Veluri et al. (2024). Figure 1 illustrates these two FD-SLM types.
@@
While e2e FD-SLMs enable more natural conversational experiences, a key challenge is that their conversational abilities degrade compared to text-based Large Language Models (TLMs), due to prolonged speech sequences and the scarcity of high-quality speech training data Zeng et al. (2024); Xie and Wu (2024). One promising solution is to transfer text abilities to speech abilities via text-speech interleaved generation, where the model first generates the assistant’s text response and uses it to guide subsequent speech generation. However, applying text guidance in e2e FD-SLMs poses significant challenges, as these models require precise time alignment between user and assistant audio streams to capture complex conversational patterns and provide timely, meaningful responses. Interleaving text tokens into the audio streams can easily disrupt this alignment, compromising the responsiveness and coherence of the generated speech.
@@
The challenges of integrating text into assistant-channel audio data can be summarized in two aspects. 1) Insertion timing. Text must be inserted precisely when the assistant begins speaking. If it is inserted too early, the assistant may respond without the full user context, causing hallucinations. If it is inserted too late, the assistant’s speech may precede the text, undermining the text guidance’s effectiveness. 2) Insertion length. The length of the inserted text is also critical. If it is too long, the assistant may overspeak and fail to adapt to subsequent user inputs; if too short, the semantic content becomes fragmented, weakening the effectiveness of text guidance Yao et al. (2025).
@@
To address these challenges, we propose a novel approach, TurnGuide, to seamlessly integrate text data into real-world double-channel spoken conversational data for the training of e2e FD-SLMs. Specifically, we dynamically segment the assistant’s speech into turns, which represent the reasonable lengths of spoken content that convey the assistant’s intended expression. Then, the assistant is trained to interleave the text-speech generation for each turn for proper guidance. By doing so, our method effectively addresses both insertion timing and insertion length challenges. Extensive experiments show that our approach significantly improves the conversational abilities of e2e FD-SLMs by not only producing semantically meaningful and coherent speech but also achieving state-of-the-art performance on various turn-taking events. To summarize our contributions:
  • We propose TurnGuide, a novel text-speech interleaved generation approach for e2e FD-SLMs grounded in natural speech data.
@@
  • We introduce a dynamic turn segmentation framework and a text-guided dialogue modeling framework within TurnGuide to address the insertion timing and length in double-channel dialogues.
  • We demonstrate through extensive experiments that TurnGuide substantially improves not only FD-SLMs’ ability to produce semantically meaningful, coherent speech but also the behaviors on various turn-taking events.
@@ 2.1 Full Duplex Speech Language Models
FD-SLMs process double-channel conversational data to interact with users seamlessly and can be categorized into cascaded and end-to-end (e2e) approaches.
2.1.1 Cascaded FD-SLMs
@@
The main focus of cascaded FD-SLMs is on defining states. For instance, LSLM Ma et al. (2025) uses states like “interrupted” and “keep speaking”, where interruptions are triggered by specific keywords. Mini-Omni 2 Xie and Wu (2024) specifies “start speaking” and “stop speaking”, halting generation upon hearing “Stop Omni!”. Freeze-Omni Wang et al. (2024) bases states on whether the user has finished speaking, prompting the model to wait or interrupt. Minmo Chen et al. (2025) simplifies its states to whether the model should speak up or keep listening. All these methods rely on synthetic conversational data to manage state transitions, limiting their ability to foster natural, human-like spoken interactions.
2.1.2 End-to-end FD-SLMs
E2e FD-SLMs, on the other hand, employ a transformer-based foundation model architecture Vaswani et al. (2017); Achiam et al. (2023) to learn from real-world double-channel conversational audio data. This e2e approach enables SLMs to autonomously capture intricate conversational dynamics from authentic human interactions, enabling more natural conversations.
The main focus of e2e FD-SLMs is on effectively modeling double-channel audio data and capturing the interactions between these channels. dGSLM Nguyen et al. (2023) introduces a Dialogue Language Model that employs distinct transformers to process each audio channel separately, using cross-attention mechanisms to model their interactions. SyncLLM Veluri et al. (2024) takes a different approach by jointly modeling the double-channel audio within a single LLM, interleaving the two audio sequences into a unified sequence. Moshi Défossez et al. (2024) adopts an RQ-Transformer architecture Lee et al. (2022), designed to predict the assistant channel’s output based on the user channel’s input at each time step. NTPP Wang et al. (2025) proposes a Next-Token-Pair Prediction training objective, enabling simultaneous prediction of audio tokens for both channels in each step.
@@ 2.2 Text-Guided Speech Generation in SLMs
SLMs enable e2e spoken interactions by generating output speech based on user speech input. However, studies have shown that SLMs exhibit significantly lower performance compared to TLMs counterparts Cui et al. (2025); Chen et al. (2025); Long et al. (2025). To mitigate this, researchers focus on integrating text to guide the speech generation process, aiming to enhance SLMs by transferring capabilities from TLMs.
Transferring abilities from TLMs to SLMs involves two key processes: representation alignment and text-speech interleaved generation. Representation alignment ensures the connection between text and speech representations within the SLM backbone, enabling knowledge transfer from text modality to speech modality. This alignment is commonly achieved through Automatic Speech Recognition (ASR) and Text-To-Speech (TTS) tasks during SLM pre-training Zeng et al. (2024); Wang et al. (2024); Chen et al. (2025). Some approaches also partially replace speech data with its corresponding text in the pre-training data Nguyen et al. (2025); Zeng et al. (2024). Subsequently, SLMs are trained to answer spoken queries by interleaving the generation of textual output with speech output, thereby effectively leveraging text information to enhance the quality of spoken responses (Zeng et al., 2024; Défossez et al., 2024).
The exploration of the text-guided speech generation for e2e FD-SLMs remains in its infancy. Currently, the only notable study in this area is Moshi Défossez et al. (2024), which proposes adding text tokens corresponding to their respective speech sounds as they occur. However, this method fragments text semantics excessively and does not significantly improve SLM performance. In contrast, our approach aggregates the text semantics and inserts text by turns, achieving substantial enhancements in SLM performance.
@@ 3 Methodology
Figure 2: Illustration of the TurnGuide approach. The first part illustrates the multi-modal turn segmentation and alignment framework, which dynamically segments assistant speech into turns and aligns the text with speech turns. The second part shows the text-guided full duplex dialogue modeling framework with two interleaving strategies.
We present the comprehensive methodology of TurnGuide, designed to enhance e2e FD-SLMs by utilizing text guidance. This approach is divided into two key frameworks: one for dynamically segmenting spoken dialogue into turns and performing speech-text data alignment, and the other for modeling text-guided full-duplex dialogue.
3.1 Dynamic Turn Segmentation and Alignment Framework
@@
As shown in the first part of Figure 2, the framework starts by applying Voice Activity Detection (VAD) to identify speech segments within the audio stream. Let $A$ represent the input audio sequence. VAD produces a list of speech segments:
@@
where each tuple $(\tau_{i}^{start},\tau_{i}^{end})$ represents the start and end timestamps of the $i$-th detected speech segment. We use the pyannote library Bredin et al. (2020) as the VAD module. Considering the VAD result is too fragmented, we merge nearby speech segments into Inter-Pausal Units (IPUs) Nguyen et al. (2023), which are defined by consecutive segments having pauses within a threshold $\tau_{IPU}=0.5$ seconds, i.e.,
@@
Enabling real-time text guidance requires the precise alignment of text and speech data. We use an Automatic Speech Recognition (ASR) module with word-level timestamp functionality to acquire the time-aligned transcript. While some speech dialogue datasets already provide time-aligned text data, our approach ensures adaptability to unlabeled dialogue datasets. The ASR process is formulated as
@@
where $(\delta_{i}^{start},\delta_{i}^{end})$ represents the start and end timestamps of word $w_{i}$. In practice, we employ the Whisper medium model Radford et al. (2023) for both ASR and timestamp extraction tasks. To enhance the stability of the timestamp extraction process, we utilize the whisper-timestamped package Louradour (2023).
@@
Finally, we align and group each word into its associated IPU using $\tau^{start}$. Additionally, we incorporate a tolerance parameter $\tau_{tol}=0.6$ seconds to adjust the word segmentation boundary slightly. This adjustment accounts for instances where, in practice, the first word in an IPU might incorrectly align with the preceding IPU due to inconsistencies in timestamping. The process is defined as
$W_{j}=\{w_{i}:\tau_{IPU_{j}}^{start}-\tau_{tol}\leq\delta_{i}^{start}\leq\tau_{IPU_{j+1}}^{start}-\tau_{tol}\},$ (5)
@@
where $W_{j}$ represents words grouped into turn $j$, and $\tau_{IPU_{j}}^{start}$ and $\tau_{IPU_{j+1}}^{start}$ represent the start times of IPU $j$ and $j+1$. We denote the speech segment corresponding to IPU $j$ as $S_{j}$. The final output consists of aligned text-speech pairs $\{(W_{j},S_{j})\}$ where each text segment is temporally aligned with its corresponding speech segment, serving as training data for text-guided full-duplex dialogue modeling.
3.2 Text-Guided Full Duplex Dialogue Modeling Framework